In a game of pool, a cue ball with an initial velocity of 1.80 m/s directed to the right (along the +x-axis) collides with the five ball, which has equal mass and is initially at rest. The five ball then moves at 1.56 m/s at an angle of 30.0° above the +x direction.
(a) Find the cue ball's final velocity.
(b) magnitude
(c) ___________ direction below the positive x-axis
Answer #1

A)

Using law of conservation of momentum in vertical direction,

1.56 sin(30) + vy = 0

Vy = - 0.78 m/s

Using law of conservation of momentum in horizontal direction,

1.8 = 1.56 cos(30) + Vx

Vx = 0.45 m/s

Velocity, v = sqrt(Vx^2 + Vy^2) = 0.90 m/s

B)

Direction, theta = arctan (Vy/Vx) = 60.02° (below the x axis)

Comment in case any doubt please rate my answer...